OpenStudy (anonymous):

solve the system by elimination -2x + 2y + 3z = 0

OpenStudy (anonymous):

UGH HELP

OpenStudy (anonymous):

I dont see a system of equations. I only see one equation.

OpenStudy (anonymous):

-2x + 2y + 3z = 0 -2x - y + z = -3 2x + 3y + 3z = 5

OpenStudy (anonymous):

show me steps i really want to understand.

OpenStudy (anonymous):

THANKS GUYS THANKS

OpenStudy (darkbluechocobo):

Alright Sorry I had to look up a video :p Give me a break

OpenStudy (darkbluechocobo):

so basically you look at it like a regular adding and subtracting equation

OpenStudy (darkbluechocobo):

and every time you see like you do in this equation -2 and +2 they are an additive inverse which means they cancel out

OpenStudy (darkbluechocobo):

so for you first step you look at the first column

OpenStudy (darkbluechocobo):

you have a -2x and a positive 2x so they cancel out so right now you have -2x left

OpenStudy (darkbluechocobo):

for the next column you don't have any additive inverses so basically you just have to solve that column

OpenStudy (darkbluechocobo):

which it would end up being 4y

OpenStudy (darkbluechocobo):

so right now you have -2x+4y

OpenStudy (darkbluechocobo):

and for the last column after adding everything it is 7z

OpenStudy (darkbluechocobo):

so for the whole thing it would go -2x+4y+7z=2

OpenStudy (darkbluechocobo):

Does that help at all?

OpenStudy (anonymous):

hahahaha i was just kidding with ya, and OHHHH i solve them then add them all together? and if they have a neg i multiply the whole thing by a positive right? and if it doesnt i just solve the equation?

OpenStudy (darkbluechocobo):

No You just solve them as they are and if it happens there is 2 of the same numbers and one is negative and the other is positive they cancel each other out

OpenStudy (darkbluechocobo):

Like in the first part you had the negative 2 and the positive 2 those canceled out just leaving another -2

OpenStudy (darkbluechocobo):

you are allowed to have negatives so you dont have to multiply by positives

OpenStudy (anonymous):

but why would i have a -2 left, what about the 3z?

OpenStudy (darkbluechocobo):

no no I am looking at this

OpenStudy (darkbluechocobo):

-2x + 2y + 3z = 0 -2x - y + z = -3 2x + 3y + 3z = 5

OpenStudy (anonymous):

on the first one

OpenStudy (darkbluechocobo):

you take them by sections

OpenStudy (darkbluechocobo):

You take the whole thing every equation

OpenStudy (darkbluechocobo):

You see how they are organized so nicely into x's y's and z's

OpenStudy (anonymous):

but on the first one i just have -2 after canceling each other out, what to i do with the 3z=o?

OpenStudy (darkbluechocobo):

See the thing is you are not looking at it as a whole with all three equations

OpenStudy (darkbluechocobo):

Look at the whole thing

OpenStudy (darkbluechocobo):

I am talking about the -2x and the +2x

OpenStudy (darkbluechocobo):

look at it vertically

OpenStudy (anonymous):

oh i solve for x then y then z?

OpenStudy (darkbluechocobo):

Exactly you get what each one equals and combine them all into one equation

OpenStudy (darkbluechocobo):

so the first would be -2x because +2x and the other -2x cancel eachother out

OpenStudy (texaschic101):

-2x + 2y + 3z = 0 -2x - y + z = -3 -->(-1) ----------------- -2x + 2y + 3z = 0 2x + y - z = 3 (result of multiplying by -1) -----------------add 0 + 3y + 2z = 3 3y + 2z = 3 we have eliminated the x in the first two equations. Now we have to eliminate x in the last two equations. -2x - y + z = -3 2x + 3y + 3z = 5 ----------------add 0 + 2y + 4z = 2 2y + 4z = 2 Now take the equations that you eliminated the x in and eliminate another variable. 3y + 2z = 3 -->(-2) 2y + 4z = 2 --------------- -6y - 4z = -6 (result of multiplying by -2) 2y + 4z = 2 ----------------add -4y + 0 = -4 -4y = -4 y = -4/-4 y = 1 now sub 1 in for y in an equation that the x was eliminated from 3y + 2z = 3 3(1) + 2z = 3 3 + 2z = 3 2z = 3 - 3 2z = 0 z = 0/2 which = 0 now that we have y and z, sub in 1 for y and 0 for z in one of the original equations -2x + 2y + 3z = 0 -2x + 2(1) + 3(0) = 0 -2x + 2 + 0 = 0 -2x + 2 = 0 -2x = -2 x = -2/-2 x = 1 now we can check our answers ( I recommend doing this) by subbing in known values into one of the original equations -2x - y + z = -3 -2(1) - 1 + 0 = -3 -2 - 1 + 0 = -3 -3 + 0 = -3 -3 = -3 (correct) If it comes out equal, then the answers are correct. x = 1, y = 1, and z = 0

OpenStudy (darkbluechocobo):

O_O

OpenStudy (anonymous):

guys you both helped me so much thank you

OpenStudy (darkbluechocobo):

o_o that paragraph is holy

OpenStudy (texaschic101):

did I explain it enough or do you have questions ?

OpenStudy (anonymous):

yeah i think i get it now and i can do it. thank you both for your time :))

OpenStudy (texaschic101):

no problem...glad to help